# Audio Tagging On An Embedded Hardware Platform

### v1

@@ 5 Experimental Analysis
Speech 50 0.69$\pm$0.17 100% 0.51$\pm$0.21 99.7% 0.75$\pm$0.12 100%
60 0.72$\pm$0.15 100% 0.52$\pm$0.21 100% 0.77$\pm$0.12 100%
70 0.73$\pm$0.14 100% 0.54$\pm$0.23 99.7% 0.78$\pm$0.12 100%
Baby cry 50 0.15$\pm$0.15 70.8% 0.20$\pm$0.16 62.7% 0.30$\pm$0.24 93%
60 0.22$\pm$0.19 82.8% 0.19$\pm$0.15 64.5% 0.34$\pm$0.25 95%
70 0.32$\pm$0.25 84.0% 0.23$\pm$0.18 66.4% 0.37$\pm$0.26 95%
Water 50 0.13$\pm$0.10 46.5% 0.21$\pm$0.16 81.9% 0.28$\pm$0.21 85%
60 0.16$\pm$0.15 60.1% 0.33$\pm$0.21 85.3% 0.28$\pm$0.22 83%
70 0.22$\pm$0.16 58.0% 0.32$\pm$0.20 80.2% 0.17$\pm$0.20 62%
Fire alarm 50 0.11$\pm$0.08 70.4% 0.42$\pm$0.21 95.9% 0.32$\pm$0.25 92%
60 0.15$\pm$0.09 68.9% 0.55$\pm$0.21 97.5% 0.39$\pm$0.27 92%
70 0.22$\pm$0.13 71.2% 0.68$\pm$0.21 98.5% 0.45$\pm$0.27 95%
Music 50 0.49$\pm$0.26 93.7% 0.56$\pm$0.23 98.0% 0.67$\pm$0.23 100%
60 0.55$\pm$0.26 96.9% 0.64$\pm$0.22 99.1% 0.72$\pm$0.24 100%
70 0.61$\pm$0.24 97.8% 0.71$\pm$0.21 99.5% 0.76$\pm$0.25 100%
@@
Hardware v/s software performance: The PC with recorded audio (scenario (3)) shows an improvement of at least 5 percentage points in confidence and a minimum of 3 percentage points improvement in Top-7 predictions compared to both the Raspberry Pi-based system (scenario (1)) and the Yeti microphone-based PC system (scenario (2))across all audio classes and audio volume levels except for “Water” class at 70dB.
@@
Impact of playback volume: Generally, the performance obtained using different systems increase with an increase in volume of the audio signals for all sound classes except for“Water” sounds. We perceptually find that increasing the volume for “Water” makes the audio recording more noisy. We also confirm using Audacity audio software [25] that the “Water” sound at 70 dB saturates resulting in audio clipping that may cause the performance degradation compared to the lower volume sounds.
@@
Impact of different microphones: We find that the AIY microphone outperforms Logitech Yeti in the detection of “Baby cry”, and “Speech” sounds for majority of the sound volume levels. It is worth mentioning that the AIY microphone is specifically designed for spoken voice detection [24], which may explain why it is able to detect speech and baby cry sounds better than that of the Logitech Mic. On the other hand, the Logitech Yeti microphone performs better than Google AIY microphone for other sounds.
@@
Performance of Audio tagging system: The audio tagging system using PANNs model on Raspberry Pi or PC shows significantly better performance in detecting “Music” and “Speech” sounds compared to other sounds. This might be due to the pre-trained PANNs model which is trained using more speech and music audio class examples compared to that of the other sounds [21].
6 Discussion: Development Issues
@@
Installation in ARM-based CPUs: Deploying the PANNs-code on a Raspberry Pi with the Raspbian operating system is not a straightforward process compared to that on a conventional computer. On the Raspberry Pi, specific library versions need to be used, and in some cases, they need to be compiled specifically for the ARM architecture of the Raspberry Pi. We provide a GitHub repository stating a detailed step-by-step guide on how to install the demonstration correctly on Raspberry Pi to promote the easier usability and easier deployment of the hardware-based AI frameworks.
@@
Temperature and latency of Raspberry Pi: We find that the temperature of the CPU in Raspberry Pi increases when we run the AI algorithms on it. To prevent overheating, the Raspberry Pi has an automatic CPU clock control mechanism that slows down the process of Raspberry Pi whenever the temperature gets too high, resulting in increased inference latency.
@@
Figure 3 shows the the temperature and latency in the Raspberry Pi when PANNs model runs for few minutes. We observe that the there is a rise in temperature of more than ${}^{\circ}$ after 14 minutes of continuous operation. After 8 minutes of running the PANNs model, the CPU temperature stabilises around ${}^{\circ}$. Meanwhile, as the temperature increases, the average inference latency also increases from approximately 0.5s to 0.6s. To mitigate this effect, it is important to incorporate heat sinks or ventilation system of the Raspberry Pi hardware container to provide sufficient cooling.
@@
Figure 3: Temperature and inference latency over time when audio tagging sytem is running on Raspberry Pi.
7 Conclusions and future work
This paper presents a deployment of PANNs-based audio tagging framework on a Raspberry Pi hardware system to stimulate further exploration of the role of hardware quality in AI applications and to promote the development of practical sound recognition technologies. We anlayse how the performance of the audio tagging software changes when deployed on hardware. We also analyse the role of the quality of microphones and audio signal volume on performance. We observe that effective device temperature management is important to maintain optimal performance and reduce inference latency. Attaining high accuracy is a balancing act involving microphone type, audio signal volume, and the chosen embedded system.
@@ 7 Conclusions and future work
Our findings suggest that the performance degrades when AI models are deployed on Raspberry Pi hardware compared to that of the software (PC) based performance. We also find that selection of an appropriate microphone is an important factor in recognising audio events. The volume levels of the audio events may also limit the performance of the AI models on Raspberry Pi. Therefore, AI models should be more robust to handle such variations due to audio volume which can be correlated with the distance of the audio source or microphone quality.
@@
In the future, we would like to explore more experiments covering a wider variety of audio event classes to understand impact of real-world versus recorded sound sources on performance. Also, we are interested in identifying optimal volume range and scale for edge device operation, measuring energy consumption to understand performance-power trade-offs and designing appropriate heat control measures to increase efficiency of the hardware-based edge devices. In terms of software, we would like to design efficient AI models to reduce computations for faster inference by exploring techniques like pruning or knowledge distillation.
8 ACKNOWLEDGEMENTS
@@
This work was supported by Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound (AI4S)”. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising. We would like to thank Dr Tim Brookes, from the Institute of Sound Recording (IoSR) at the University of Surrey, for his invaluable advice and for providing the volume meter Precision Gold IEC 651.

### v2

@@ 4.3.  Performance metrics
[…]
5.1.  Performance differences between the embedded system and the software system
@@
The PC with recorded audio (scenario (3)) shows an improvement of at least 5 percentage points in confidence and a minimum of 3 percentage points improvement in Top-7 predictions compared to both the Raspberry Pi-based system (scenario (1)) and the Yeti microphone-based PC system (scenario (2)) across all audio classes and audio volume levels except for “Water” class at 70dB.
5.2.  Impact of playback volume
@@
Generally, the performance obtained using different systems increase with an increase in volume of the audio signals for all sound classes except for “Water” sounds. We perceptually find that increasing the volume for “Water” makes the audio recording more noisy. We also confirm using Audacity audio software [25] that the “Water” sound at 70 dB saturates resulting in audio clipping that may cause the performance to degrade compared to the lower volume sounds.
5.3.  Impact of different microphones
@@
We find that the AIY microphone outperforms the Logitech Yeti in the detection of “Baby cry”, and “Speech” sounds for majority of the sound volume levels. It is worth mentioning that the AIY microphone is specifically designed for spoken voice detection [24], which may explain why it is able to detect speech and baby cry sounds better than that of the Logitech Mic. On the other hand, the Logitech Yeti microphone performs better than the Google AIY microphone for other sounds.
5.4.  Performance of Audio tagging system
@@
The audio tagging system using the PANNs model on the Raspberry Pi or PC shows significantly better performance in detecting “Music” and “Speech” sounds compared to other sounds. This might be due to the pre-trained PANNs model which is trained using more speech and music audio class examples compared to that of the other sounds [21].
.  iscussion: Development Issues
6.1.  Installation in ARM-based CPUs
@@
Deploying the PANNs-code on a Raspberry Pi with the Raspbian operating system is not a straightforward process compared to that on a conventional computer. On the Raspberry Pi, specific library versions need to be used, and in some cases, they need to be compiled specifically for the ARM architecture of the Raspberry Pi. We provide a GitHub repository [26] giving a detailed step-by-step guide on how to install the demonstration correctly on Raspberry Pi to promote the easier usability and easier deployment of the hardware-based AI frameworks.
6.2.  Temperature and latency of Raspberry Pi
@@
We find that the temperature of the Raspberry Pi’s CPU increases when we run the AI algorithms on it. To prevent overheating, the Raspberry Pi has an automatic CPU clock control mechanism that slows down the processing on the Raspberry Pi whenever the temperature gets too high, resulting in increased inference latency.
@@
Figure 3 shows the the temperature and latency in the Raspberry Pi when the PANNs model runs for few minutes. We observe that the there is a rise in temperature of more than 25C after 14 minutes of continuous operation. After 8 minutes of running the PANNs model, the CPU temperature stabilises around 79C. Meanwhile, as the temperature increases, the average inference latency also increases from approximately 0.5s to 0.6s. To mitigate this effect, it is important to incorporate heat sinks or a ventilation system for the Raspberry Pi hardware container to provide sufficient cooling.
@@
Figure 3: Temperature and inference latency over time when audio tagging system is running on Raspberry Pi.
.  onclusions and future work
@@
In this paper, we presented a deployment of a PANNs-based audio tagging framework on a Raspberry Pi hardware system. Our findings suggest that the performance degrades when AI models are deployed on a Raspberry Pi compared to software-based (PC) performance. We also find that the selection of an appropriate microphone is an important factor in recognizing audio events, and the volume levels of the audio events may limit the performance of the AI models.
We observe that effective device temperature management is important to maintain optimal performance and reduce inference latency. When exploring heat control methods like fans, it is important to consider the negative impact they may have on the audio estimates made by the device. Fans generate noise, known as egonoise, which can reduce the quality of the audio captured by the microphone. Attaining high accuracy is a balancing act involving microphone type, audio signal volume, and the chosen embedded system.
@@
In the future, we would like to explore more experiments covering a wider variety of audio event classes to understand the impact of real-world versus recorded sound sources on performance. Analyzing different neural network architectures beyond PANNs would also be valuable. Moreover, we are interested in identifying the range of audio volumes suitable for optimal operation of these edge devices, determining the minimum and maximum levels within which the system can operate reliably, and designing appropriate heat control measures to increase the efficiency of hardware-based edge devices.
Furthermore, future studies could investigate techniques to optimize models for edge devices, such as pruning or knowledge distillation, to reduce computations and enable faster inference. Measuring the energy consumption of the Raspberry Pi during inference would provide insights into performance-power trade-offs. Considering a broader range of edge devices beyond the Raspberry Pi, such as microcontrollers and FPGAs, would provide a more comprehensive understanding of the challenges and opportunities in deploying AI models on resource-constrained devices.
In conclusion, our work serves as a starting point for investigation into the deployment of audio tagging models on edge devices. Despite the limitations, we believe our findings will help guide the next directions of research in this area. By addressing the identified challenges and exploring the suggested future directions, we can strengthen the applicability and impact of audio tagging models in real-world edge computing scenarios.
cknowledgements
@@
This work was supported by Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound (AI4S)”. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.